What is tge magnetic moment associated with a coil of one turn, are of a cross section 10^-4m^2 carrying current of 2A

Grade:12

1 Answers

Khimraj
7 years ago
Magnetic moment associated with coil is given by M= NIAWhere N is no of turnsI is current And A is cross section area So M = 1*2*10^-4 = 2×10^-4 amp/m²